A block of mass 1.5 kg is allowed to slide down along a quadrant of a circle from the horizontal position in reaching to the bottom its velocity is 8 M per second the work done is overcoming the friction is 12j the radius of circle is

Grade:11

1 Answers

Arun
5 years ago
Using the conservation of energy

mgh=1/2 mv
²+ 12

given mass =1.5Kg

velocity=8m/s

1.5*10*h=1/2*1.5*8
²+12

15h=60

h=4m

so the radius also should be equal to 4m